# QServe: W4A8KV4 Quantization and System Co-design for Efficient LLM Serving

### v2

@@ Abstract
Quantization can accelerate large language model (LLM) inference. Going beyond INT8 quantization, the research community is actively exploring even lower precision, such as INT4. Nonetheless, state-of-the-art INT4 quantization techniques only accelerate low-batch, edge LLM inference, failing to deliver performance gains in large-batch, cloud-based LLM serving. We uncover a critical issue: existing INT4 quantization methods suffer from significant runtime overhead (20-90%) when dequantizing either weights or partial sums on GPUs. To address this challenge, we introduce QoQ, a W4A8KV4 quantization algorithm with 4-bit weight, 8-bit activation, and 4-bit KV cache. QoQ stands for quattuor-octō-quattuor, which represents 4-8-4 in Latin. QoQ is implemented by the QServe inference library that achieves measured speedup. The key insight driving QServe is that the efficiency of LLM serving on GPUs is critically influenced by operations on low-throughput CUDA cores. Building upon this insight, in QoQ algorithm, we introduce progressive quantization that can allow low dequantization overhead in W4A8 GEMM. Additionally, we develop SmoothAttention to effectively mitigate the accuracy degradation incurred by 4-bit KV quantization. In the QServe system, we perform compute-aware weight reordering and take advantage of register-level parallelism to reduce dequantization latency. We also make fused attention memory-bound, harnessing the performance gain brought by KV4 quantization. As a result, QServe improves the maximum achievable serving throughput of Llama-3-8B by 1.2$\times$ on A100, 1.4$\times$ on L40S; and Qwen1.5-72B by 2.4$\times$ on A100, 3.5$\times$ on L40S, compared to TensorRT-LLM. Remarkably, QServe on L40S GPU can achieve even higher throughput than TensorRT-LLM on A100. Thus, QServe effectively reduces the dollar cost of LLM serving by 3$\times$. Code is released at https://github.com/mit-han-lab/qserve.
footnotetext: *: The first three authors contribute equally to this project and are listed in the alphabetical order. Yujun Lin leads the quantization algorithm, Haotian Tang and Shang Yang lead the GPU kernels and the serving system.
I Introduction
@@
Figure 1: QServe achieves higher throughput when running Llama models on L40S compared with TensorRT-LLM on A100, effectively saves the dollar cost for LLM serving by 3$\times$ through system-algorithm codesign. See Table IV for absolute throughput numbers and precision choices in TensorRT-LLM.
Large language models (LLMs) have demonstrated remarkable capability across a broad spectrum of tasks, exerting a profound influence on our daily lives.
@@
However, the colossal size of LLMs makes their deployment extremely challenging, necessitating the adoption of quantization techniques for efficient inference. State-of-the-art integer quantization algorithms can be divided into three categories: 8-bit weight and 8-bit activation (W8A8), 4-bit weight and 16-bit activation (W4A16), 4-bit weight 4-bit activation (W4A4) quantization. The former two methods are considered nearly lossless in terms of accuracy. In contrast, W4A4 quantization introduces a notable accuracy degradation, although it is anticipated to offer superior throughput in return by mapping its computations onto high-throughput 4-bit tensor cores. Unfortunately, this anticipated performance boost has not been consistently observed across current GPU platforms. For instance, the state-of-the-art W4A4 serving system, Atom [44], exhibits 20-25% lower performance than its W4A16 and W8A8 counterpart in TensorRT-LLM when running the Llama-2-7B [34] model on A100 GPUs. That said, the research community has yet to find a precision combination superior to W4A16 and W8A8 for efficient cloud LLM serving.
@@
In this paper, we reveal a critical observation: current 4-bit integer quantization methods experience significant overhead, ranging from 20% to 90%, during the dequantization of weights or partial sums on current-generation GPUs. For example, W4A16 quantization performs computation on FP16 tensor cores while the weights are in INT4, so weight dequantization is required in the GEMM kernel. On the other hand, for W4A4 quantization, to achieve reasonable accuracy, W4A4 methods must apply per-group quantization to both weights and activation, sharing FP16 scaling factors on a sub-channel basis. For example, the state-of-the-art W4A4 quantization method, QuaRot [2], reports a significant 0.2 perplexity degradation after switching from per-group quantization to per-channel quantization. This per-group quantization design requires an integer to floating-point dequantization for partial sums (since INT4 tensor cores produce INT32 partial sums), which operates on the slower CUDA cores within the sequential main loop of W4A4 GEMM. On data center GPUs like A100, a CUDA core operation is as expensive as 50 INT4 tensor core operations.
@@
Therefore, reducing overhead on CUDA cores is crucial for achieving optimal throughput in LLM serving. Guided by this principle, we introduce QoQ (Quattuor-Octō-Quattuor, or 4-8-4 in Latin) algorithm which quantizes LLMs to W4A8KV4 precision: 4-bit weights, 8-bit activations and 4-bit KV caches. Additionally, we present QServe, which provides efficient system support for W4A8KV4 quantization.
@@
In the QoQ algorithm, we introduce progressive group quantization. This method first quantizes weights to 8 bits using per-channel FP16 scales, then quantizes these 8-bit intermediates to 4 bits. This approach ensures that all GEMMs are performed on INT8 tensor cores. Additionally, we mitigate accuracy loss from KV4 quantization through SmoothAttention, which shifts the challenge of activation quantization from keys to queries, the latter of which are not quantized.
@@
We evaluate seven widely-used LLMs using QServe on A100 and L40S GPUs, and compare their maximum achievable throughput against state-of-the-art systems, including TensorRT-LLM (in FP16, W8A8, and W4A16 configurations), Atom [44] (in W4A4), and QuaRot [2] (in W4A4). On A100 GPUs, QServe achieves 1.2-2.4$\times$ higher throughput over the best-performing configuration of TensorRT-LLM, and 2.5-2.9$\times$ higher throughput compared to Atom and QuaRot. On L40S GPUs, QServe records an even more significant 1.5-3.5$\times$ throughput improvement over TensorRT-LLM. Notably, we manage to accommodate the same batch size on the L40S while consistently achieving higher serving throughput than TensorRT-LLM on A100 for six of the eight models tested, thereby significantly reducing LLM serving cost by 3$\times$.
II Background
II-A Large Language Models
@@
In attention blocks, $\mathbf{x}$ first undergoes linear projection to obtain $\mathbf{q}\in\mathbb{R}^{N\times HD},\mathbf{k},\mathbf{v}\in\mathbb{R}^{N\times H_{KV}D}$, where $H_{KV}$ is the number of key/value heads. We have $H=H_{KV}$ in the standard multi-head attention (MHA), while recent methods [34, 17, 18] also employ grouped-query attention (GQA) [1] with $H=rH_{KV}(r\in\mathbb{Z})$. We concatenate $\mathbf{k},\mathbf{v}$ with pre-computed KV cache features of $S$ previous tokens to obtain $\mathbf{K},\mathbf{V}\in\mathbb{R}^{(S+N)\times H_{KV}D}$ and compute attention using:
$\small\mathbf{o}_{h}=\text{softmax}\left(\frac{\mathbf{q}_{h}\mathbf{K}_{h_{KV}}^{T}}{\sqrt{D}}\right)\mathbf{V}_{h_{KV}},\quad h_{KV}=\left\lfloor\frac{h}{r}\right\rfloor.$ (1)
@@ II-B Integer Quantization
$\small\mathbf{Q}_{\mathbf{X}}=\left\lceil\frac{\mathbf{X}}{s}+z\right\rfloor,s=\frac{\mathbf{X}_{\max}-\mathbf{X}_{\min}}{q_{\max}-q_{\min}},z=\left\lceil q_{\min}-\frac{\mathbf{X}_{\min}}{s}\right\rfloor,$ (2)
@@
$\small\hat{\mathbf{X}}=Q\left(\mathbf{X}\right)=\left(\mathbf{Q}_{\mathbf{X}}-z\right)\cdot s$ (3)
@@ III Motivation
III-A W4A8KV4 Has Superior Roofline Over W8A8, W4A16
@@
III-B Why Not W4A4KV4: Main Loop Overhead in GEMM
@@
A natural follow-up question would be: “Why do we not choose the even more aggressive W4A4?” W4A4 starts to achieve better theoretical GEMM performance when $m$, the number of input sequences, exceeds 78, as 4-bit tensor cores are twice as performant compared to their 8-bit counterparts. However, apart from the significant accuracy degradation, which will be discussed in Section VI, we demonstrate that such theoretical performance gains cannot be realized on existing GPU architectures (Ampere and Hopper). As in Figure 2b, existing W4A4 serving systems Atom [44] and QuaRot [2] are even significantly slower than the W16A16 solution from TensorRT-LLM.
@@ IV QoQ Quantization
IV-A Progressive Group Quantization
Figure 6: Progressive Group Quantization first employs per-channel INT8 quantization with protective range [-119, 119], followed by per-group INT4 quantization, so that the dequantized intermediate values remain within the INT8 range for computation. Bottom: prior methods directly applies per-group INT4 quantization on weights, followed by per-channel INT8 quantization on scale factors. Thus the dequantized intermediate values may exceed the INT8 range, necessitating further dequantization to floating-point values for computation.
To enhance the accuracy of low-bit quantization, group quantization is commonly utilized [44, 23, 12]. However, as outlined in Section 5, the dequantization overhead in the system implementation can negate these accuracy improvements. To tackle this issue, we introduce progressive group quantization, as depicted in Figure 6.
@@ IV-A Progressive Group Quantization
$\small\hat{\mathbf{W}}={\mathbf{Q}_{\mathbf{W}}}^{(0)}_{\mathrm{s8}}\cdot\mathbf{s}^{(0)}_{\mathrm{fp16}},$ (4)
@@
$\small{{\mathbf{Q}}_{\mathbf{W}}}_{\mathrm{s8}}^{(0)}=\left({\mathbf{Q}_{\mathbf{W}}}_{\mathrm{u4}}-\mathbf{z}_{\mathrm{u4}}\right)\cdot\mathbf{s}^{(1)}_{\mathrm{u8}},$ (5)
@@
Protective Quantization Range
naïvely applying Equation 4 and 5 does not guarantee that the intermediate dequantized weights perfectly lie in the 8-bit integer representation range. For example, after INT8 quantization, a group of 8-bit weights lie in $[-113,120]$. 4-bit asymmetric quantization will yield a scale factor of (120113)/(150)=16$\lceil$(120--113)/(15-0$\rfloor$=16⌈ ( 120 - - 113 ) / ( 15 - 0 ) ⌋ = 16 and a zero point of 0113/16=7$\lceil$ 0--113/16$\rfloor$=7⌈ 0 - - 113 / 16 ⌋ = 7. Thus value 120 is quantized into 120/16+7=15$\lceil$ 120/16+7$\rfloor$=15⌈ 120 / 16 + 7 ⌋ = 15. It will be dequantized into $(15-7)*16=128$ which is beyond the max 8-bit integer 127. One straightforward solution is to turn on the saturation option in the arithmetic instructions during dequantization. However, simply applying saturation will severely damage the computation throughput, reducing speed by as much as 67%.
@@
$\small\hat{q}_{s8}=\lfloor\frac{{q}_{s8}}{{s}_{u8}}\rceil\cdot{{s}_{u8}}\leq{q}_{s8}+\frac{1}{2}{{s}_{u8}}.$
@@
$\small\hat{q}_{s8}\leq 127\rightarrow{q}_{s8}\leq 127-\frac{1}{2}{{s}_{u8}}\rightarrow{q}_{s8}\leq 119.5$
@@
Compared to previous two-level quantization
progressive group quantization introduces two levels of scales $\mathbf{s}^{(0)}_{\mathrm{fp16}}$ and $\mathbf{s}^{(1)}_{\mathrm{u8}}$. Prior studies such as VSQuant and DoubleQuant in QLoRA [9] also introduce two levels of scales to reduce the memory footprint of group-wise scaling factors. In contrast to our quantization flow, previous approaches directly apply group quantization with the target precision and then perform per-channel quantization on the group-wise floating-point scaling factors, as shown in the bottom of Figure 6:
$\small\hat{\mathbf{W}}={\mathbf{Q}_{\mathbf{W}}}_{\mathrm{s4}}\cdot\mathbf{s}_{\mathrm{fp16}},\;\;\;\hat{\mathbf{s}}_{\mathrm{fp16}}={\mathbf{s}}^{(1)}_{\mathrm{u8}}\cdot\mathbf{s}^{(0)}_{\mathrm{fp16}}$ (6)
@@
DGQ [43] also follows the quantization scheme of VSQuant and DoubleQuant, but enforces restrictions on scaling factors to make sure that all computation can be mapped onto INT8 tensor cores. However, the DGQ serving system separates dequantization kernel with the GEMM kernel. Consequently, the end-to-end latency of W4A8 GEMM in DGQ is even slower than the W8A8 GEMM in cuBLAS, failing to demonstrate the memory bandwidth advantage of 4-bit weight quantization. In contrast, our QoQ introduces a protective range, allowing us to fuse dequantization operations into the W4A8 GEMM kernel with full register-level parallelism, minimizing CUDA core overhead. Thus, our QServe’s W4A8 per-group GEMM achieves 1.5$\times$ speedup over the W8A8 cuBLAS GEMM.
IV-B SmoothAttention

### v3

@@ Abstract
Quantization can accelerate large language model (LLM) inference. Going beyond INT8 quantization, the research community is actively exploring even lower precision, such as INT4. Nonetheless, state-of-the-art INT4 quantization techniques only accelerate low-batch, edge LLM inference, failing to deliver performance gains in large-batch, cloud-based LLM serving. We uncover a critical issue: existing INT4 quantization methods suffer from significant runtime overhead (20-90%) when dequantizing either weights or partial sums on GPUs. To address this challenge, we introduce QoQ, a W4A8KV4 quantization algorithm with 4-bit weight, 8-bit activation, and 4-bit KV cache. QoQ stands for quattuor-octō-quattuor, which represents 4-8-4 in Latin. QoQ is implemented by the QServe inference library that achieves measured speedup. The key insight driving QServe is that the efficiency of LLM serving on GPUs is critically influenced by operations on low-throughput CUDA cores. Building upon this insight, in QoQ algorithm, we introduce progressive quantization that can allow low dequantization overhead in W4A8 GEMM. Additionally, we develop SmoothAttention to effectively mitigate the accuracy degradation incurred by 4-bit KV quantization. In the QServe system, we perform compute-aware weight reordering and take advantage of register-level parallelism to reduce dequantization latency. We also transfer theoretical memory saving brought by KV4 attention into measured speedup using QServe. As a result, QServe improves the maximum achievable serving throughput of Llama-3-8B by 1.2$\times$ on A100, 1.4$\times$ on L40S; and Qwen1.5-72B by 2.4$\times$ on A100, 3.5$\times$ on L40S, compared to TensorRT-LLM. Remarkably, QServe on L40S GPU can achieve even higher throughput than TensorRT-LLM on A100. Code is released at https://github.com/mit-han-lab/omniserve.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
@@
Figure 1: QServe achieves higher throughput when running Llama models on L40S compared with TensorRT-LLM on A100, effectively saves the dollar cost for LLM serving by 3$\times$ through system-algorithm codesign. See Table 4 for absolute throughput numbers and precision choices in TensorRT-LLM.
@@
Large language models (LLMs) have demonstrated remarkable capability across a broad spectrum of tasks, exerting a profound influence on our daily lives. However, the colossal size of LLMs makes their deployment extremely challenging, necessitating the adoption of quantization techniques for efficient inference. State-of-the-art integer quantization algorithms can be divided into three categories: 8-bit weight and 8-bit activation (W8A8), 4-bit weight and 16-bit activation (W4A16), 4-bit weight 4-bit activation (W4A4) quantization. The former two methods are considered nearly lossless in terms of accuracy. In contrast, W4A4 quantization introduces a notable accuracy degradation, although it is anticipated to offer superior throughput in return by mapping its computations onto high-throughput 4-bit tensor cores. Unfortunately, this anticipated performance boost has not been consistently observed across current GPU platforms. For instance, the state-of-the-art W4A4 serving system, Atom Zhao et al. (2023), exhibits 20-25% lower performance than its W4A16 and W8A8 counterpart in TensorRT-LLM when running the Llama-2-7B Touvron et al. (2023b) model on A100 GPUs. That said, the research community has yet to find a precision combination superior to W4A16 and W8A8 for efficient cloud LLM serving.
@@
In this paper, we reveal a critical observation: current 4-bit integer quantization methods experience significant overhead, ranging from 20% to 90%, during the dequantization of weights or partial sums on current-generation GPUs. For example, W4A16 quantization performs computation on FP16 tensor cores while the weights are in INT4, so weight dequantization is required in the GEMM kernel. On the other hand, for W4A4 quantization, to achieve reasonable accuracy, W4A4 methods must apply per-group quantization to both weights and activation, sharing FP16 scaling factors on a sub-channel basis. For example, the state-of-the-art W4A4 quantization method, QuaRot Ashkboos et al. (2024), reports a significant 0.2 perplexity degradation after switching from per-group quantization to per-channel quantization. This per-group quantization design requires an integer to floating-point dequantization for partial sums (since INT4 tensor cores produce INT32 partial sums), which operates on the slower CUDA cores within the sequential main loop of W4A4 GEMM. On data center GPUs like A100, a CUDA core operation is as expensive as 50 INT4 tensor core operations. Therefore, reducing bit precision not necessarily speeds up LLM inference.
@@
To achieve optimal LLM serving throughput, We introduce QoQ (Quattuor-Octō-Quattuor, or 4-8-4 in Latin) algorithm which quantizes LLMs to W4A8KV4 precision: 4-bit weights, 8-bit activations and 4-bit KV caches. Additionally, we present QServe, which provides efficient system support for W4A8KV4 quantization.
@@
In the QoQ algorithm, we introduce progressive group quantization. This method first quantizes weights to 8 bits using per-channel FP16 scales with a protective range of $[-119,119]$, then quantizes these 8-bit intermediates to 4 bits. This approach ensures that all GEMMs are performed on INT8 tensor cores. Additionally, we mitigate accuracy loss from KV4 quantization through SmoothAttention, which shifts the challenge of activation quantization from keys to queries, the latter of which are not quantized.
@@
We evaluate seven widely-used LLMs using QServe on A100 and L40S GPUs, and compare their maximum achievable throughput against state-of-the-art systems, including TensorRT-LLM (in FP16, W8A8, and W4A16 configurations), Atom Zhao et al. (2023) (in W4A4), and QuaRot Ashkboos et al. (2024) (in W4A4). On A100 GPUs, QServe achieves 1.2-2.4$\times$ higher throughput over the best-performing configuration of TensorRT-LLM, and 2.5-2.9$\times$ higher throughput compared to Atom and QuaRot. On L40S GPUs, QServe records an even more significant 1.5-3.5$\times$ throughput improvement over TensorRT-LLM. Notably, we manage to accommodate the same batch size on the L40S while consistently achieving higher serving throughput than TensorRT-LLM on A100 which is 3$\times$ more expensive for six of the eight models tested.
2 Background
2.1 Large Language Models
@@
In attention blocks, $\mathbf{x}$ first undergoes linear projection to obtain $\mathbf{q}\in\mathbb{R}^{N\times HD},\mathbf{k},\mathbf{v}\in\mathbb{R}^{N\times H_{KV}D}$, where $H_{KV}$ is the number of key/value heads. We have $H=H_{KV}$ in the standard multi-head attention (MHA), while recent methods Touvron et al. (2023b); Jiang et al. (2023; 2024) also employ grouped-query attention (GQA) Ainslie et al. (2023) with $H=rH_{KV}(r\in\mathbb{Z})$. We concatenate $\mathbf{k},\mathbf{v}$ with pre-computed KV cache features of $S$ previous tokens to obtain $\mathbf{K},\mathbf{V}\in\mathbb{R}^{(S+N)\times H_{KV}D}$ and compute attention using:
$\mathbf{o}_{h}=\text{softmax}\left(\frac{\mathbf{q}_{h}\mathbf{K}_{h_{KV}}^{T}}{\sqrt{D}}\right)\mathbf{V}_{h_{KV}},\quad h_{KV}=\left\lfloor\frac{h}{r}\right\rfloor$ (1)
@@ 2.2 Integer Quantization
$\mathbf{Q}_{\mathbf{X}}=\left\lceil\frac{\mathbf{X}}{s}+z\right\rfloor,s=\frac{\mathbf{X}_{\max}-\mathbf{X}_{\min}}{q_{\max}-q_{\min}},z=\left\lceil q_{\min}-\frac{\mathbf{X}_{\min}}{s}\right\rfloor$
@@
$\hat{\mathbf{X}}=Q\left(\mathbf{X}\right)=\left(\mathbf{Q}_{\mathbf{X}}-z\right)\cdot s$ (3)
@@ 3 Motivation
3.1 W4A8KV4 Has Superior Roofline Over W8A8, W4A16
@@
3.2 Why Not W4A4KV4: Main Loop Overhead in GEMM
@@
A natural follow-up question would be: “Why do we not choose the even more aggressive W4A4?” W4A4 starts to achieve better theoretical GEMM performance when $m$, the number of input sequences, exceeds 78, as 4-bit tensor cores are twice as performant compared to their 8-bit counterparts. However, apart from the significant accuracy degradation, which will be discussed in Section 6, we demonstrate that such theoretical performance gains cannot be realized on existing GPU architectures (Ampere and Hopper). As in Figure 2b, existing W4A4 serving systems Atom Zhao et al. (2023) and QuaRot Ashkboos et al. (2024) are even significantly slower than the W16A16 solution from TensorRT-LLM.
@@ 4 QoQ Quantization
4.1 Progressive Group Quantization
Figure 6: Progressive Group Quantization first employs per-channel INT8 quantization with protective range [-119, 119], followed by per-group INT4 quantization, so that the dequantized intermediate values remain within the INT8 range for computation.
To enhance the accuracy of low-bit quantization, group quantization is commonly utilized Zhao et al. (2023); Lin et al. (2024); Frantar et al. (2022). However, as outlined in Section 5, the dequantization overhead in the system implementation can negate these accuracy improvements. To tackle this issue, we introduce progressive group quantization, as depicted in Figure 6.
@@ 4.1 Progressive Group Quantization
$\hat{\mathbf{W}}={\mathbf{Q}_{\mathbf{W}}}^{(0)}_{\mathrm{s8}}\cdot\mathbf{s}^{(0)}_{\mathrm{fp16}},$ (4)
@@
${{\mathbf{Q}}_{\mathbf{W}}}_{\mathrm{s8}}^{(0)}=\left({\mathbf{Q}_{\mathbf{W}}}_{\mathrm{u4}}-\mathbf{z}_{\mathrm{u4}}\right)\cdot\mathbf{s}^{(1)}_{\mathrm{u8}},$ (5)
@@
Protective Quantization Range.
Naively applying Equation 4 and 5 does not guarantee that the intermediate dequantized weights perfectly lie in the 8-bit integer representation range (i.e., $[-128,127]$). For example, after INT8 quantization, a group of 8-bit weights lie in $[-113,120]$. 4-bit asymmetric quantization will yield a scale factor of (120113)/(150)=16$\lceil$(120--113)/(15-0$\rfloor$=16⌈ ( 120 - - 113 ) / ( 15 - 0 ) ⌋ = 16 and a zero point of 0113/16=7$\lceil$ 0--113/16$\rfloor$=7⌈ 0 - - 113 / 16 ⌋ = 7. Thus value 120 is quantized into 120/16+7=15$\lceil$ 120/16+7$\rfloor$=15⌈ 120 / 16 + 7 ⌋ = 15. It will be dequantized into $(15-7)*16=128$ which is beyond the max 8-bit integer 127. One straightforward solution is to turn on the saturation option in the arithmetic instructions during dequantization. However, simply applying saturation will severely damage the computation throughput, reducing speed by as much as 67%.
@@
$\hat{q}_{s8}=\lfloor\frac{{q}_{s8}}{{s}_{u8}}\rceil\cdot{{s}_{u8}}\leq{q}_{s8}+\frac{1}{2}{{s}_{u8}}$
@@
$\hat{q}_{s8}\leq 127\rightarrow{q}_{s8}\leq 127-\frac{1}{2}{{s}_{u8}}\rightarrow{q}_{s8}\leq 119.5$
@@
Compared to previous two-level quantization.
Progressive group quantization introduces two levels of scales $\mathbf{s}^{(0)}_{\mathrm{fp16}}$ and $\mathbf{s}^{(1)}_{\mathrm{u8}}$. Prior studies such as VSQuant and DoubleQuant in QLoRA Dettmers et al. (2023a) also introduce two levels of scales to reduce the memory footprint of group-wise scaling factors. In contrast to our quantization flow, previous approaches directly apply group quantization with the target precision and then perform per-channel quantization on the group-wise floating-point scaling factors, as shown in the bottom of Figure 6:
$\hat{\mathbf{W}}={\mathbf{Q}_{\mathbf{W}}}_{\mathrm{s4}}\cdot\mathbf{s}_{\mathrm{fp16}},\;\;\;\hat{\mathbf{s}}_{\mathrm{fp16}}={\mathbf{s}}^{(1)}_{\mathrm{u8}}\cdot\mathbf{s}^{(0)}_{\mathrm{fp16}}$ (6)
@@
DGQ Zhang et al. (2023) also follows the quantization scheme of VSQuant and DoubleQuant, but enforces restrictions on scaling factors to make sure that all computation can be mapped onto INT8 tensor cores. However, the DGQ serving system separates dequantization kernel with the GEMM kernel. Consequently, the end-to-end latency of W4A8 GEMM in DGQ is even slower than the W8A8 GEMM in cuBLAS, failing to demonstrate the memory bandwidth advantage of 4-bit weight quantization. In contrast, our QoQ introduces a protective range, allowing us to fuse dequantization operations into the W4A8 GEMM kernel with full register-level parallelism, minimizing CUDA core overhead. Thus, our QServe’s W4A8 per-group GEMM achieves 1.5$\times$ speedup over the W8A8 cuBLAS GEMM.
4.2 SmoothAttention